FROM A CIRCULAR DISC OF RADIUS R AND MASS 9M, A SMALL DISC OF RADIUS R/3 IS REMOVED FROM THE DISC. THE M.O.INERTIA OF THE REMAINING DISC ABOUT AN AXIS PERPENDICULAR TO THE PLANE OF THE DISC AND PASSING THRO' IT'S CENTER IS: 

Grade:11

3 Answers

Ritvik Gautam
13 years ago

it will be the same...

Shahid Akbar
10 years ago
Its will be 4MR*2 BY using parallel axis theorem
Jay Gohil
6 years ago
Let I1 be the moment of Inertia of bigger circle and I2 of smaller circleI=I1- I2=9mr^2/2-mr^2/18=80mr^2/18=40mr^2/9